Given the sequence 7, 14, 28, 56, ..., which expression shown would give the tenth term?
Answer by checkley77(12844)

L=A*R^(N-1)
L IS THE LAST NUMBER.
A=THE FIRST NUMBER.
R IS THE RATIO.
N IS THE Nth TERM.
L=7*2^(10-1)
L=7*2^9
L=7*512
L=3,584 IS THE 10th TERM IN THE SEQUENCE.